The two identical charges have a magnitude of ±1.00×10-6 Coulombs which separated by 1 cm. What is the magnitude of the force separating the charges?


1
Expert's answer
2021-12-20T10:32:10-0500

"F=k\frac{q_1 q_2}{r^2}=9\cdot10^{9}\cdot\frac{1\cdot10^{-6}\cdot1\cdot10^{-6}}{0.01^2}=90\ (N)"